OpenStudy (anonymous):

The function f(x)=104.5x^2 - 1501.5x + 6016, models the death rate per year per 100,000 males, f(x), for men who average x hours of sleep each night. How many hours of sleep, to the nearest tenth of an hour, corresponds to the minimum death rate, to the nearest whole number?

OpenStudy (anonymous):

f'(x)=209x-1501.5; f'(x)=0 when x= 1501.5/209=7.18. So, the minimum death rate occurs with about 7 hours of sleep.

OpenStudy (anonymous):

how can we solve this without using calculus?

OpenStudy (anonymous):

The axis of symmetry of a parabola is y=-b/2a. In this case, that would be y=1501.1/209=7.2, which rounds to 7 hours.

OpenStudy (anonymous):

Aren't we looking for x here, since the question is 'how many hours of sleep'? [argh, i really don't understand the problem.:( ]

OpenStudy (anonymous):

Sorry, you're right. It should be x=-b/2a. But the answer's the same.